OpenStudy (anonymous):

Write the expression using exponents: 4•4•4•4•4•4•a•a•a•a. A. 4a10 B. 4^6a^4 C. 6a^4 D. 6^4a^4

OpenStudy (cggurumanjunath):

b

OpenStudy (anonymous):

I am leaning to B! But not sure.

OpenStudy (cggurumanjunath):

4*4=4^2 a*a*a=a^3